# High-dimensional covariance estimation by pairwise likelihood truncation

### v1

@@ 5 Conclusion and final remarks
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$. If $\Pr(\hat{\mathcal{E}}\subseteq{\mathcal{E}})\to 1$ as $n\to\infty$, then $\hat{\lambda}=O(\gamma)O_{p}(m_{0})$.
@@
By lemma assumption, $\Pr\{\hat{\mathcal{E}}\subseteq{\mathcal{E}}\}\to 1$ as $n\to\infty$. Thus, it suffices to consider the case where $\hat{\mathcal{E}}\subseteq{\mathcal{E}}$. By KKT condition (28) and the definition of $\hat{\lambda}$, the next pair $jk$ ($j<k$) about to enter $\hat{\mathcal{E}}$ satisfies $nS_{jk}^{2}\leq\gamma(S_{jk}^{2}+S_{jj}S_{kk})$ and
@@
Since $m_{0}$ is fixed, the optimizer of the former converges to the latter, i.e. $\hat{\mathbf{w}}_{\xi}\to{\mathbf{w}}_{\xi}^{\ast}$ in probability where ${\mathbf{w}}_{\xi}^{\ast}$ is such that ${\mathbf{J}}_{\xi}{\mathbf{w}}_{\xi}^{\ast}={\mathbf{h}}_{\xi}$. By Condition C3, all elements of ${\mathbf{w}}_{\xi}^{\ast}$ are nonzero, implying the same for $\hat{\mathbf{w}}_{\xi}$. Similar to the proof in Lemma 5, elements of $\hat{\mathbf{w}}_{\setminus\xi}$ corresponding to the marginal scores are all equal one. ∎

### v2

@@ 5 Conclusion and final remarks
Under Conditions C1-C4, for any constant $r_{1}\in(0,\frac{r}{8+r})$ and $r_{2}=\min(\frac{r}{32+4r},\\ \frac{4+r}{32+4r}-\frac{r_{1}}{4})$, let $m_{0}=o(n^{r_{2}-c})$ for some arbitrarily small $c\in(0,\min\{r_{1},r_{2}\})$. If $\Pr(\hat{\mathcal{E}}\subseteq{\mathcal{E}})\to 1$ as $n\to\infty$, then $\hat{\lambda}=O(\gamma)O_{p}(m_{0})$.
@@
By the assumption of the Lemma, $\Pr\{\hat{\mathcal{E}}\subseteq{\mathcal{E}}\}\to 1$ as $n\to\infty$. Thus, it suffices to consider the case where $\hat{\mathcal{E}}\subseteq{\mathcal{E}}$. By KKT condition (28) and the definition of $\hat{\lambda}$, the next pair $jk$ ($j<k$) about to enter $\hat{\mathcal{E}}$ satisfies $nS_{jk}^{2}\leq\gamma(S_{jk}^{2}+S_{jj}S_{kk})$ and
@@
Since $m_{0}$ is fixed, the optimizer of the former converges to the latter, i.e., $\hat{\mathbf{w}}_{\xi}\to{\mathbf{w}}_{\xi}^{\ast}$ in probability where ${\mathbf{w}}_{\xi}^{\ast}$ is such that ${\mathbf{J}}_{\xi}{\mathbf{w}}_{\xi}^{\ast}={\mathbf{h}}_{\xi}$. By Condition C3, all elements of ${\mathbf{w}}_{\xi}^{\ast}$ are nonzero, implying the same for $\hat{\mathbf{w}}_{\xi}$. Similar to the proof in Lemma 5, elements of $\hat{\mathbf{w}}_{\setminus\xi}$ corresponding to the marginal scores are all equal one. ∎